OpenStudy (anonymous):

Find the 30th term of the following sequence 1,7,13,19,...

OpenStudy (zarkon):

\[a_n=1+(n-1)\cdot 6\]

OpenStudy (anonymous):

Now you get to solve it :)

OpenStudy (zarkon):

I already know how...you try.

OpenStudy (anonymous):

What do the vaiables stand for? where do I plug in certain numbers?

OpenStudy (zarkon):

\[a_{30}=1+(30-1)\cdot6=1+29\cdot6\]